Grade 11Mechanics

An automobile on a road moves with a speed of 54km/h.The radius of its wheels is 0.35m.What is the average torque transmitted by its brakes to a wheel,if the vehicle is brought to rest in 15s?The moment of inertia of the wheel about its axis of rotation is 3kgm2.

Apu
7 Years agoGrade 11

1 Answer

Arun
7 Years ago
Hello Dear.
Given Conditions ⇒
Initial linear velocity(u) = 54 km/hr.
 = 54 × 5/18 [Changing into m/s.]
 = 15 m/s.
Radius of the wheel(r) = 0.35 m.
Now, Using the Formula,
Initial Angular Velocity (ω₁) = u/r
ω₁ = 15/0.35
ω₁ = 42.86 radian/sec.
Also, Final Angular Velocity (ω₂) = 0
Now, Using the Formula,
Angular Acceleration (α) = (ω₂ - ω₁)/t
 = (0 - 42.86)/15
 = - 2.86 radian/sec²
Moment of Inertia about the axis of the rotation(I) = 3 kg-m²
∴ Torque = I × α
⇒ Torque = 3 × -2.86
⇒ Torque = -8.58 Nm.

Hence, the average negative torque is -8.58 N-m.
Regards
Arun (askIITiams forum expert)